OpenStudy (anonymous):

Choose the correct slope of the line shown on the graph below. graph of a line passing through points negative 2 comma 1 and 0 comma negative 3 2 −2 one half −one half

OpenStudy (anonymous):

I think -2?

OpenStudy (e.mccormick):

Do you know the formula for finding slope? Ah, an answer... let me check.

OpenStudy (e.mccormick):

Yah.

OpenStudy (anonymous):

No i dont really know the formula.

OpenStudy (e.mccormick):

Oh, well, the formula might help. You did get it right, but let me do the formula.

OpenStudy (anonymous):

so your points can be simplified as - (-2, 1) and (0,-3) If there are two points (x1,y1) and (x2,y2) slope is given By = M = y2-y1/x2-x1.

OpenStudy (anonymous):

u hence get = -3-1/0-(-2) which is -4/+2 which is -2.

OpenStudy (e.mccormick):

\[m=\frac{y_1-y_2}{x_1-x_2}\]Where m is the slope, and the x and y parts are from the two points. So you are given \((x_1,y_1)\) and \((x_2,y_2)\) or pick them off a graph, and then put them into the formula.

OpenStudy (anonymous):

thanks both :)

OpenStudy (e.mccormick):

So if you get that formula down, you can use it for any slope. They also call it: \[\frac{\mathrm{Change\, in\, y}}{\mathrm{Change\, in\, x}}=\frac{\Delta y}{\Delta x}\]